OpenStudy (anonymous):

Find the difference of (-5x^2 - 4x^3 - 3) - (9x^3 + 3x - 8x^2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2. Find the difference of (-5x^2 - 4x^3 - 3) - (9x^3 + 3x - 8x^2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2. @Mathematics

OpenStudy (anonymous):

−5x^2 − 4x^3 − 3) − (9x^3 + 3x − 8x^2) = −5x^2 − 4x^3 − 3 − 9x^3 − 3x + 8x^2 Combine like terms: = (−4 − 9)x^3 + (−5 + 8)x^2 − 3x − 3 = −13x^3 + 3x^2 − 3x − 3

OpenStudy (anonymous):

Medal plz :$